OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Write your answer as a mixed number. Base 3 1/2 feet, altitude 3/4 feet.

OpenStudy (anonymous):

Area of a parallelogram is Base*Height, so you just need to multiply those numbers: \[(3\frac{1}{2})(\frac{3}{4}) = (\frac{7}{2})(\frac{3}{4}) = \frac{21}{8} = 2\frac{5}{8}\]